### v2

@@ Appendix C Baseline
In Appendix C, we compare PILOT against six representative zero-shot anomaly detection methods, summarized as follows:
  • SAA [Cao et al.(2023b)Cao, Xu, Sun, Cheng, Du, Gao, and Shen]: Relies on a pretrained Grounding DINO to propose candidate anomaly regions, which are then segmented by the Segment Anything Model (SAM) without any additional training.
  • CoCoOp [Zhou et al.(2022a)Zhou, Yang, Loy, and Liu]: Replaces fixed CLIP prompt tokens with learnable context vectors while keeping the backbone frozen. Offers both a shared “unified” prompt and class-specific prompts to adapt flexibly to unseen categories.
  • WinCLIP [Jeong et al.(2023)Jeong, Zou, Kim, Zhang, Ravichandran, and Dabeer]: Splits the image into overlapping windows, encodes each via CLIP to produce localized anomaly scores, and aggregates window predictions through an ensemble of token embeddings—improving spatial localization without fine-tuning.
  • APRIL-GAN [Chen et al.(2023)Chen, Han, and Zhang]: Builds on WinCLIP by refining text prompts for tighter vision–language alignment and injecting lightweight learnable projections into the visual encoder, thereby sharpening anomaly masks while retaining zero-shot generality.
  • AnomalyCLIP [Zhou et al.(2024)Zhou, Pang, Tian, He, and Chen]: Learns object-agnostic “normal” and “anomalous” text prompts, then employs a dense patch-to-prompt attention module (DPAM) to align each visual patch embedding with the appropriate prompt for pixel-level localization.
  • AdaCLIP [Cao et al.(2024)Cao, Zhang, Frittoli, Cheng, Shen, and Boracchi]: Combines static prompts shared across all inputs for coarse adaptation with dynamic, image-conditioned prompts for fine-grained adjustment. Both prompt sets are jointly learned on auxiliary anomaly data to balance detection accuracy and localization precision.
@@ D.2 Anomaly State Descriptors
Let $\boldsymbol{\kappa}=(\kappa_{1},\dots,\kappa_{D})\in\mathbb{R}^{D}$ be a vector of input scores, where each $\kappa_{i}\in\mathbb{R}$ denotes the unnormalized score for the $i$-th dimension of interest, and write the probability simplex as
@@
where $\Delta^{D}$ is the set of all $D$-dimensional probability distributions.
@@
where the scalar $\tau(\boldsymbol{\kappa})\in\mathbb{R}$ is chosen so that $\sum_{i=1}^{D}\chi_{i}=1$. Intuitively, any entry with $\kappa_{i}\leq\tau(\boldsymbol{\kappa})$ is set to zero, and the remaining values are shifted, squared, and renormalized. This yields a differentiable, piecewise-quadratic mapping that retains only the most salient dimensions in its support. In PILOT, injecting Gumbel noise before applying entmax1.5 ensures that only the highest-scoring attribute cues survive to influence the final text embedding, providing robust, sparse weighting during training.
@@ E.1 Comparison with SOTA full-shot method
Industrial Image-level Anomaly Detection (AUROC, AP)
@@
DTD-Synthetic (97.7, 98.4) (97.5, 97.3) (96.8, 96.9) (96.2, 96.3)
Industrial Pixel-level Anomaly Localization (AUROC, AUPRO)
@@
DTD-Synthetic (98.0, 90.2) (98.2, 98.1) (98.3, 98.5) (97.6, 97.7)
In this comparison, we evaluate PILOT alongside three state-of-the-art fully supervised methods—PatchCore [Roth et al.(2022)Roth, Pemula, Zepeda, Schölkopf, Brox, and Gehler], CDO [Cao et al.(2023a)Cao, Xu, Liu, and Shen], and RD4AD [Deng and Li(2022)]—across six widely used industrial benchmarks. Here, a full-shot method refers to a model trained with full access to samples from the target domain. While these baselines leverage direct access to target domain knowledge, PILOT operates solely with pretrained vision–language representations and requires no information from the target domain. Remarkably, PILOT achieves detection accuracy on par with these full-shot models, even when faced with unseen textures or novel defect types. It matches PatchCore’s memory-based robustness, rivals the collaborative feature learning of CDO, and maintains high sensitivity to both global and local irregularities, comparable to RD4AD’s reconstruction-driven approach. However, there remains room for improvement in pixel-level localization accuracy, which we identify as a promising direction for future research. These results underscore the significant potential of zero-shot anomaly detection frameworks in real-world industrial scenarios.
@@ E.2 Pseudo label Statistic
Table 6: TTA Dataset selection distribution statistics across datasets under different pseudo-labeling ratios $\rho$. M1: Shannon Evenness, M2: Coefficient of Variation, M3: Percentage of Noisy Label.
Dataset Class $\rho=$5% $\rho=$15% $\rho=$25% $\rho=$35% $\rho=$50% M1 M2 M3 M1 M2 M3 M1 M2 M3 M1 M2 M3 M1 M2 M3 MVTec AD 15 0.76 1.21 10.3% 0.97 0.45 8.1% 0.93 0.71 7.8% 0.95 0.59 11.2% 1.00 0.00 27.3% VisA 12 0.79 0.97 12.0% 0.98 0.27 13.2% 0.98 0.35 14.2% 0.99 0.31 16.1% 1.00 0.00 31.4% BTAD 3 0.79 0.76 21.6% 0.86 0.70 12.6% 0.93 0.50 17.3% 0.95 0.39 20.4% 1.00 0.00 36.5% SDD 1 0.72 0.59 0.0% 0.61 0.69 7.6% 0.63 0.68 6.3% 0.74 0.58 10.1% 1.00 0.00 39.6% DAGM 10 0.73 1.35 3.9% 0.88 0.76 1.3% 0.95 0.56 0.7% 0.97 0.43 0.9% 1.00 0.00 6.9% DTD-Synthetic 12 0.72 1.34 3.0% 0.96 0.47 6.1% 0.95 0.53 4.6% 0.96 0.48 5.2% 1.00 0.00 20.7%
In Table 6 we report three statistics that describe how our TTA selection (See Section 2.5 of the main paper) distributes pseudo-labels over the target dataset $D_{\mathrm{tar}}$. Let $B=\lvert D_{\mathrm{tar}}\rvert$ and recall that we form the top-$\rho$ set $A$ and bottom-$\rho$ set $N$, with $S=A\cup N$. For each class $i$ in the set of classes $\mathcal{I}$, we write $g_{i}$ for the fraction of samples in $S$ belonging to class $i$.
@@
    where $\mu(g)$ and $\sigma(g)$ are the mean and standard deviation of $\{g_{i}\}$. A low $\mathrm{CV}$ signals a balanced class-wise count. This is referred as M2 in Table 6.
  • Noisy label rate is the fraction of indices in $S$ whose pseudo-label $\tilde{y}_{j}$ disagrees with the ground-truth label $y_{j}$. This is referred as M3 in Table 6.
As $\rho$ grows, we observe that $E$ increases and $\mathrm{CV}$ decreases, confirming that selecting a larger fraction of samples both diversifies class coverage and balances the counts. At the same time, the noisy label rate rises, reflecting the classic trade-off: higher coverage under TTA comes at the expense of pseudo-label accuracy.
@@ E.3 Other TTA Method
To better understand the impact of different TTA strategies on anomaly detection performance, we conduct an ablation study comparing several representative TTA methods. As shown in Table 7, we evaluate TENT [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell], BN Adaptation [Lim et al.(2023)Lim, Kim, Choo, and Choi], and soft pseudo labeling alongside our proposed PILOT framework.
@@
Table 7 presents the results of different TTA strategies on industrial anomaly detection benchmarks. Both TENT and BN Adaptation exhibit weak performance in this setting, particularly for pixel-level localization, suggesting that these conventional TTA methods are not well suited for adapting CLIP-based models. Recent literature has also observed that methods such as TENT, which adapt only batch normalization parameters, often struggle when applied to models with large distribution shifts or strong vision-language priors, as is the case with CLIP [Liang et al.(2025)Liang, He, and Tan]. In contrast, the soft pseudo labeling strategy leads to more stable improvements, although its cautious approach may be less aggressive in leveraging adaptation signals. Our findings suggest that confidence-based pseudo-labeling, as used in the PILOT TTA framework, provides the most effective adaptation and yields the best overall performance in this scenario.
@@ E.7 Prompt Aggregation Ablation
Table 11: Ablation for prompt aggregation. “w/o TTA” and “w/ TTA” indicate performance without and with TTA, respectively. “Extra” indicates whether additional trainable parameters are used. OP: orthogonal projector as used in PILOT; GF: Gate Fusion [Arevalo et al.(2017)Arevalo, Solorio, Montes-y Gómez, and González]; CA: Cross Attention [Chen et al.(2021)Chen, Fan, and Panda].
@@
Table 11 compares three prompt aggregation strategies and reveals that methods like Gate Fusion [Arevalo et al.(2017)Arevalo, Solorio, Montes-y Gómez, and González] and Cross Attention [Chen et al.(2021)Chen, Fan, and Panda], while capable of improving image-level anomaly detection, tend to destabilize pixel-level localization during TTA. While these approaches may leverage the attribute memory bank as a semantic anchor, their extra parameters are directly updated using noisy pseudo-labels and only image-level supervision. This makes them prone to parameter drift, which can erode the stability needed for precise localization under domain shift. In contrast, the orthogonal projector used in PILOT achieves reliable gains across both detection and localization tasks without introducing extra parameters, as it keeps the attribute embedding fixed as a semantic reference and allows only the prompt representation to adapt. This encourages adaptation in directions that are truly complementary, while avoiding redundancy and overfitting.
@@ 5 $\circ$ damaged, paint peeling, weld defect
Having established the inherent diversity within our learnable prompt pool $\mathcal{P}$, we further investigated how this diversity corresponds to meaningful anomaly characteristics expressed in natural language. To accomplish this, we conducted a nearest-neighbor analysis using FAISS [Douze et al.(2024)Douze, Guzhva, Deng, Johnson, Szilvasy, Mazaré, Lomeli, Hosseini, and Jégou], comparing anomaly prompt embeddings ($T_{a}$) to embeddings generated from textual templates of the form “a photo of a [anomaly state] object.” Various anomaly state descriptors were used to represent potential real-world anomalies.
@@ Appendix G Medical anomaly localization results
Table 14: Comparison of model performance on medical anomaly localization datasets. Pixel-level results (AUROC, AUPRO) in parentheses. Best entries are marked with bold and second-best entries with underline.
Medical Pixel-level Anomaly Localization (AUROC, AUPRO) Dataset SAA CoCoOp WinCLIP APRIL-GAN AnomalyCLIP AdaCLIP PILOT w/o TTA ISIC (83.8, 74.2) (51.7, 15.9) (83.3, 55.1) (89.4, 77.2) (87.4, 74.5) (84.4, 54.5) (87.8, 74.2) CVC-ColonDB (71.8, 31.5) (40.5, 2.6) (70.3, 32.5) (78.4, 64.6) (88.5, 79.3) (88.0, 63.9) (88.6, 78.2) CVC-ClinicDB (66.2, 29.1) (34.8, 2.4) (51.2, 13.8) (80.5, 60.7) (93.0, 81.6) (94.4, 73.5) (92.1, 78.5) Kvasir (45.9, 13.3) (44.1, 3.5) (69.7, 24.5) (75.0, 36.2) (93.2, 59.9) (94.6, 26.2) (94.0, 58.8) Mean (66.9, 37.0) (42.8, 6.1) (68.6, 31.5) (80.8, 59.7) (90.5, 73.8) (90.4, 54.5) (90.6, 72.4)
Medical anomaly localization results are reported in Table 14, with all evaluations conducted without TTA. Since medical datasets lack explicit “normal” examples, our TTA strategy cannot be applied; nevertheless, PILOT still produces competitive localization maps, achieving the highest mean AUROC and the second-best AUPRO across benchmarks. These results underscore the generality of our dual-branch prompt learning framework: even in the absence of adaptation, dynamically weighting prompts with the most relevant anomaly cues deliver state-of-the-art performance in medical domains.
@@ Appendix I Compared with AnovL
AnovL [Deng et al.(2023)Deng, Zhang, Bao, and Li] is, to our knowledge, the only existing test-time adaptation method tailored for zero-shot anomaly localization. It introduces a lightweight adapter on the visual token stream and refines pixel-level anomaly maps by minimizing a self-supervised reconstruction loss; the adapter learns to reconstruct features of normal patches, so that anomalous regions, whose reconstructions incur higher error, become more salient.
@@
By design, AnovL’s adaptation objective targets only pixel-level refinement and does not update any image-level scoring components. As a result, AnovL achieves modest gains in localization accuracy but has no impact on image-level detection (Table 16). In contrast, PILOT’s TTA strategy optimizes a high-confidence, image-level objective by adaptively updating the learnable prompt parameters according to their relevance for each test image, which leads to consistent improvements on both image- and pixel-level metrics. This dual benefit arises because PILOT emphasizes prompt parameters most aligned with the test data, while the fixed attribute embeddings act as semantic anchors to stabilize adaptation and prevent parameter drift.
To investigate the complementarity of these adaptation schemes, we conducted an ablation in which PILOT’s image-level TTA is combined with AnovL’s pixel reconstruction loss (row “+ AnoVL Loss”). Although this hybrid approach retains PILOT’s adaptive prompt weighting mechanism and the visual adapter architecture, it failed to improve either localization or detection. We hypothesize that the reconstruction and pseudo-label losses impose conflicting objectives on the same prompt parameters, driving contradictory update signals that destabilize adaptation and degrade performance.
Table 16: Comparison between PILOT and AnovL on two benchmarks (MVTec AD and VisA). “w/o TTA” and “w/ TTA” indicate performance without and with test-time adaptation, respectively. Arrows denote change under TTA: $\uparrow$(improvement), $\downarrow$(decline), $\circ$(no significant change, $\pm 0.5$).
@@
In Appendix K, we present high-level pseudocode for both the training and test-time adaptation procedures of PILOT. Algorithm 1 details the end-to-end optimization on the auxiliary dataset $D_{\mathrm{aux}}$, where we jointly update the learnable prompt pool $\mathcal{P}$, attribute memory bank $\mathcal{U}$, and final visual projection layer using combined image-level and pixel-level objectives. Algorithm 2 then describes our confidence-based, image-level objective TTA scheme on an unlabeled target set $D_{\mathrm{tar}}$, in which only the relevant prompt parameters and their associated $r_{k}$, $s_{k}$ embeddings are fine-tuned to align with high-confidence pseudo-labels.
@@
Figure 4: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the bottle category in the MVTec AD dataset. All results are generated by PILOT.
@@
Figure 5: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the grid category in the MVTec AD dataset. All results are generated by PILOT.
@@
Figure 6: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the metal nut category in the MVTec AD dataset. All results are generated by PILOT.
@@
Figure 7: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the wood category in the MVTec AD dataset. All results are generated by PILOT.
@@
Figure 8: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the cable category in the MVTec AD dataset. All results are generated by PILOT.
@@
Figure 9: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the chewinggum category in the VisA dataset. All results are generated by PILOT.
@@
Figure 10: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the cashew category in the VisA dataset. All results are generated by PILOT.
@@
Figure 11: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the pcb4 category in the VisA dataset. All results are generated by PILOT.
@@
Figure 12: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the pcb2 category in the VisA dataset. All results are generated by PILOT.
@@
Figure 13: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the 01 category in the BTAD dataset. All results are generated by PILOT.
@@
Figure 14: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the 02 category in the BTAD dataset. All results are generated by PILOT.
@@
Figure 15: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the 03 category in the BTAD dataset. All results are generated by PILOT.
@@
Figure 16: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the electrical commutators category in the SDD dataset. All results are generated by PILOT.
@@
Figure 17: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the class 1 category in the DAGM dataset. All results are generated by PILOT.
@@
Figure 18: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the class 2 category in the DAGM dataset. All results are generated by PILOT.
@@
Figure 19: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the class 3 category in the DAGM dataset. All results are generated by PILOT.
@@
Figure 20: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the class 4 category in the DAGM dataset. All results are generated by PILOT.
@@
Figure 21: Visualization of input images (top), ground-truth masks (middle), and model-generated anomaly maps (bottom) for the Blotchy category in the DTD dataset. All results are generated by PILOT.

### v3

@@ Appendix C Baseline
In Appendix C, we compare PILOT against six representative zero‑shot anomaly detection methods, summarized as follows:
  • SAA [saa]: Relies on a pretrained Grounding DINO to propose candidate anomaly regions, which are then segmented by the Segment Anything Model (SAM) without any additional training.
  • CoCoOp [CoCoOp]: Replaces fixed CLIP prompt tokens with learnable context vectors while keeping the backbone frozen. Offers both a shared “unified” prompt and class‑specific prompts to adapt flexibly to unseen categories.
  • WinCLIP [winclip]: Splits the image into overlapping windows, encodes each via CLIP to produce localized anomaly scores, and aggregates window predictions through an ensemble of token embeddings—improving spatial localization without fine‑tuning.
  • APRIL‑GAN [April_GAN]: Builds on WinCLIP by refining text prompts for tighter vision–language alignment and injecting lightweight learnable projections into the visual encoder, thereby sharpening anomaly masks while retaining zero‑shot generality.
  • AnomalyCLIP [AnomalyCLIP]: Learns object‑agnostic “normal” and “anomalous” text prompts, then employs a dense patch‑to‑prompt attention module (DPAM) to align each visual patch embedding with the appropriate prompt for pixel‑level localization.
  • AdaCLIP [AdaCLIP]: Combines static prompts shared across all inputs for coarse adaptation with dynamic, image‑conditioned prompts for fine‑grained adjustment. Both prompt sets are jointly learned on auxiliary anomaly data to balance detection accuracy and localization precision.
@@ D.2 Anomaly State Descriptors
Let $\boldsymbol{\kappa}=(\kappa_{1},\dots,\kappa_{D})\in\mathbb{R}^{D}$ be a vector of input scores, where each $\kappa_{i}\in\mathbb{R}$ denotes the unnormalized score for the ii‑th dimension of interest, and write the probability simplex as
@@
where $\Delta^{D}$ is the set of all DD‑dimensional probability distributions.
@@
where the scalar $\tau(\boldsymbol{\kappa})\in\mathbb{R}$ is chosen so that $\sum_{i=1}^{D}\chi_{i}=1$. Intuitively, any entry with $\kappa_{i}\leq\tau(\boldsymbol{\kappa})$ is set to zero, and the remaining values are shifted, squared, and renormalized. This yields a differentiable, piecewise‑quadratic mapping that retains only the most salient dimensions in its support. In PILOT, injecting Gumbel noise before applying entmax1.5 ensures that only the highest‑scoring attribute cues survive to influence the final text embedding, providing robust, sparse weighting during training.
@@ E.1 Comparison with SOTA full-shot method
Industrial Image‑level Anomaly Detection (AUROC, AP)
@@
DTD‑Synthetic (97.7, 98.4) (97.5, 97.3) (96.8, 96.9) (96.2, 96.3)
Industrial Pixel‑level Anomaly Localization (AUROC, AUPRO)
@@
DTD‑Synthetic (98.0, 90.2) (98.2, 98.1) (98.3, 98.5) (97.6, 97.7)
In this comparison, we evaluate PILOT alongside three state-of-the-art fully supervised methods—PatchCore [roth2022towards], CDO [cao2023collaborative], and RD4AD [deng2022anomaly]—across six widely used industrial benchmarks. Here, a full-shot method refers to a model trained with full access to samples from the target domain. While these baselines leverage direct access to target domain knowledge, PILOT operates solely with pretrained vision–language representations and requires no information from the target domain. Remarkably, PILOT achieves detection accuracy on par with these full-shot models, even when faced with unseen textures or novel defect types. It matches PatchCore’s memory-based robustness, rivals the collaborative feature learning of CDO, and maintains high sensitivity to both global and local irregularities, comparable to RD4AD’s reconstruction-driven approach. However, there remains room for improvement in pixel-level localization accuracy, which we identify as a promising direction for future research. These results underscore the significant potential of zero-shot anomaly detection frameworks in real-world industrial scenarios.
@@ E.2 Pseudo label Statistic
Table 6: TTA Dataset selection distribution statistics across datasets under different pseudo‑labeling ratios $\rho$. M1: Shannon Evenness, M2: Coefficient of Variation, M3: Percentage of Noisy Label.
Dataset Class $\rho=$5% $\rho=$15% $\rho=$25% $\rho=$35% $\rho=$50% M1 M2 M3 M1 M2 M3 M1 M2 M3 M1 M2 M3 M1 M2 M3 MVTec AD 15 0.76 1.21 10.3% 0.97 0.45 8.1% 0.93 0.71 7.8% 0.95 0.59 11.2% 1.00 0.00 27.3% VisA 12 0.79 0.97 12.0% 0.98 0.27 13.2% 0.98 0.35 14.2% 0.99 0.31 16.1% 1.00 0.00 31.4% BTAD 3 0.79 0.76 21.6% 0.86 0.70 12.6% 0.93 0.50 17.3% 0.95 0.39 20.4% 1.00 0.00 36.5% SDD 1 0.72 0.59 0.0% 0.61 0.69 7.6% 0.63 0.68 6.3% 0.74 0.58 10.1% 1.00 0.00 39.6% DAGM 10 0.73 1.35 3.9% 0.88 0.76 1.3% 0.95 0.56 0.7% 0.97 0.43 0.9% 1.00 0.00 6.9% DTD‑Synthetic 12 0.72 1.34 3.0% 0.96 0.47 6.1% 0.95 0.53 4.6% 0.96 0.48 5.2% 1.00 0.00 20.7%
[…]
@@
    where $\mu(g)$ and $\sigma(g)$ are the mean and standard deviation of $\{g_{i}\}$. A low $\mathrm{CV}$ signals a balanced class‑wise count. This is referred as M2 in Table 6.
  • Noisy label rate is the fraction of indices in $S$ whose pseudo‑label $\tilde{y}_{j}$ disagrees with the ground‑truth label $y_{j}$. This is referred as M3 in Table 6.
As $\rho$ grows, we observe that $E$ increases and $\mathrm{CV}$ decreases, confirming that selecting a larger fraction of samples both diversifies class coverage and balances the counts. At the same time, the noisy label rate rises, reflecting the classic trade‑off: higher coverage under TTA comes at the expense of pseudo‑label accuracy.
@@ E.3 Other TTA Method
To better understand the impact of different TTA strategies on anomaly detection performance, we conduct an ablation study comparing several representative TTA methods. As shown in Table 7, we evaluate TENT [wang2020tent], BN Adaptation [lim2023ttn], and soft pseudo labeling alongside our proposed PILOT framework.
@@
Table 7 presents the results of different TTA strategies on industrial anomaly detection benchmarks. Both TENT and BN Adaptation exhibit weak performance in this setting, particularly for pixel-level localization, suggesting that these conventional TTA methods are not well suited for adapting CLIP-based models. Recent literature has also observed that methods such as TENT, which adapt only batch normalization parameters, often struggle when applied to models with large distribution shifts or strong vision-language priors, as is the case with CLIP [liang2025comprehensive]. In contrast, the soft pseudo labeling strategy leads to more stable improvements, although its cautious approach may be less aggressive in leveraging adaptation signals. Our findings suggest that confidence-based pseudo-labeling, as used in the PILOT TTA framework, provides the most effective adaptation and yields the best overall performance in this scenario.
@@ E.7 Prompt Aggregation Ablation
Table 11: Ablation for prompt aggregation. “w/o TTA” and “w/ TTA” indicate performance without and with TTA, respectively. “Extra” indicates whether additional trainable parameters are used. OP: orthogonal projector as used in PILOT; GF: Gate Fusion [arevalo2017gated]; CA: Cross Attention [chen2021crossvit].
@@
Table 11 compares three prompt aggregation strategies and reveals that methods like Gate Fusion [arevalo2017gated] and Cross Attention [chen2021crossvit], while capable of improving image-level anomaly detection, tend to destabilize pixel-level localization during TTA. While these approaches may leverage the attribute memory bank as a semantic anchor, their extra parameters are directly updated using noisy pseudo-labels and only image-level supervision. This makes them prone to parameter drift, which can erode the stability needed for precise localization under domain shift. In contrast, the orthogonal projector used in PILOT achieves reliable gains across both detection and localization tasks without introducing extra parameters, as it keeps the attribute embedding fixed as a semantic reference and allows only the prompt representation to adapt. This encourages adaptation in directions that are truly complementary, while avoiding redundancy and overfitting.
@@ 5 $\circ$ damaged, paint peeling, weld defect
Having established the inherent diversity within our learnable prompt pool $\mathcal{P}$, we further investigated how this diversity corresponds to meaningful anomaly characteristics expressed in natural language. To accomplish this, we conducted a nearest-neighbor analysis using FAISS [douze2024faiss], comparing anomaly prompt embeddings ($T_{a}$) to embeddings generated from textual templates of the form “a photo of a [anomaly state] object.” Various anomaly state descriptors were used to represent potential real-world anomalies.
@@ Appendix G Medical anomaly localization results
Table 14: Comparison of model performance on medical anomaly localization datasets. Pixel‑level results (AUROC, AUPRO) in parentheses. Best entries are marked with bold and second‑best entries with underline.
Medical Pixel-level Anomaly Localization (AUROC, AUPRO) Dataset SAA CoCoOp WinCLIP APRIL-GAN AnomalyCLIP AdaCLIP PILOT w/o TTA ISIC (83.8, 74.2) (51.7, 15.9) (83.3, 55.1) (89.4, 77.2) (87.4, 74.5) (84.4, 54.5) (87.8, 74.2) CVC‑ColonDB (71.8, 31.5) (40.5, 2.6) (70.3, 32.5) (78.4, 64.6) (88.5, 79.3) (88.0, 63.9) (88.6, 78.2) CVC‑ClinicDB (66.2, 29.1) (34.8, 2.4) (51.2, 13.8) (80.5, 60.7) (93.0, 81.6) (94.4, 73.5) (92.1, 78.5) Kvasir (45.9, 13.3) (44.1, 3.5) (69.7, 24.5) (75.0, 36.2) (93.2, 59.9) (94.6, 26.2) (94.0, 58.8) Mean (66.9, 37.0) (42.8, 6.1) (68.6, 31.5) (80.8, 59.7) (90.5, 73.8) (90.4, 54.5) (90.6, 72.4)
Medical anomaly localization results are reported in Table 14, with all evaluations conducted without TTA. Since medical datasets lack explicit “normal” examples, our TTA strategy cannot be applied; nevertheless, PILOT still produces competitive localization maps, achieving the highest mean AUROC and the second‑best AUPRO across benchmarks. These results underscore the generality of our dual‑branch prompt learning framework: even in the absence of adaptation, dynamically weighting prompts with the most relevant anomaly cues deliver state‑of‑the‑art performance in medical domains.
@@ Appendix I Compared with AnovL
AnovL [deng2023anovl] is, to our knowledge, the only existing test‑time adaptation method tailored for zero‑shot anomaly localization. It introduces a lightweight adapter on the visual token stream and refines pixel‑level anomaly maps by minimizing a self‑supervised reconstruction loss; the adapter learns to reconstruct features of normal patches, so that anomalous regions, whose reconstructions incur higher error, become more salient.
@@
By design, AnovL’s adaptation objective targets only pixel‑level refinement and does not update any image‑level scoring components. As a result, AnovL achieves modest gains in localization accuracy but has no impact on image‑level detection (Table 16). In contrast, PILOT’s TTA strategy optimizes a high-confidence, image-level objective by adaptively updating the learnable prompt parameters according to their relevance for each test image, which leads to consistent improvements on both image- and pixel-level metrics. This dual benefit arises because PILOT emphasizes prompt parameters most aligned with the test data, while the fixed attribute embeddings act as semantic anchors to stabilize adaptation and prevent parameter drift.
To investigate the complementarity of these adaptation schemes, we conducted an ablation in which PILOT’s image‑level TTA is combined with AnovL’s pixel reconstruction loss (row “+ AnoVL Loss”). Although this hybrid approach retains PILOT’s adaptive prompt weighting mechanism and the visual adapter architecture, it failed to improve either localization or detection. We hypothesize that the reconstruction and pseudo‑label losses impose conflicting objectives on the same prompt parameters, driving contradictory update signals that destabilize adaptation and degrade performance.
Table 16: Comparison between PILOT and AnovL on two benchmarks (MVTec AD and VisA). “w/o TTA” and “w/ TTA” indicate performance without and with test‑time adaptation, respectively. Arrows denote change under TTA: $\uparrow$(improvement), $\downarrow$(decline), $\circ$(no significant change, $\pm 0.5$).
@@
In Appendix K, we present high‑level pseudocode for both the training and test‑time adaptation procedures of PILOT. Algorithm 1 details the end‑to‑end optimization on the auxiliary dataset $D_{\mathrm{aux}}$, where we jointly update the learnable prompt pool $\mathcal{P}$, attribute memory bank $\mathcal{U}$, and final visual projection layer using combined image‑level and pixel‑level objectives. Algorithm 2 then describes our confidence‑based, image-level objective TTA scheme on an unlabeled target set $D_{\mathrm{tar}}$, in which only the relevant prompt parameters and their associated $r_{k}$, $s_{k}$ embeddings are fine‑tuned to align with high‑confidence pseudo‑labels.
@@
Figure 4: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the bottle category in the MVTec AD dataset. All results are generated by PILOT.
@@
Figure 5: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the grid category in the MVTec AD dataset. All results are generated by PILOT.
@@
Figure 6: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the metal nut category in the MVTec AD dataset. All results are generated by PILOT.
@@
Figure 7: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the wood category in the MVTec AD dataset. All results are generated by PILOT.
@@
Figure 8: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the cable category in the MVTec AD dataset. All results are generated by PILOT.
@@
Figure 9: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the chewinggum category in the VisA dataset. All results are generated by PILOT.
@@
Figure 10: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the cashew category in the VisA dataset. All results are generated by PILOT.
@@
Figure 11: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the pcb4 category in the VisA dataset. All results are generated by PILOT.
@@
Figure 12: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the pcb2 category in the VisA dataset. All results are generated by PILOT.
@@
Figure 13: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the 01 category in the BTAD dataset. All results are generated by PILOT.
@@
Figure 14: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the 02 category in the BTAD dataset. All results are generated by PILOT.
@@
Figure 15: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the 03 category in the BTAD dataset. All results are generated by PILOT.
@@
Figure 16: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the electrical commutators category in the SDD dataset. All results are generated by PILOT.
@@
Figure 17: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the class 1 category in the DAGM dataset. All results are generated by PILOT.
@@
Figure 18: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the class 2 category in the DAGM dataset. All results are generated by PILOT.
@@
Figure 19: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the class 3 category in the DAGM dataset. All results are generated by PILOT.
@@
Figure 20: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the class 4 category in the DAGM dataset. All results are generated by PILOT.
@@
Figure 21: Visualization of input images (top), ground‐truth masks (middle), and model‐generated anomaly maps (bottom) for the Blotchy category in the DTD dataset. All results are generated by PILOT.